A political candidate has asked you to conduct a poll to determine what percentage of people support her.
If the candidate only wants a 8% margin of error at a 90% confidence level, what size of sample is needed?
Give your answer in whole people.

Answer by Boreal(15235)

.08=z*SE
0.08=1.645*sqrt((.5*.5)/n), which is the most conservative estimation of the proportion.
sqrtn)=1.645*sqrt(.5*.5)/0.08=1.645*.5/0.08=10.28
n=105.70 or 106 people